# A Context-Driven Approach for Co-Auditing Smart Contracts with The Support of GPT-4 code interpreter

### v1

@@ A. Affirming the Efficacy of Our CCL-Based Chunking Approach
B. Adequacy of Our CAQ-Prompt in Practical Application Our meticulously designed CAQ prompt is structured around four core components: judgment of vulnerabilities, explaining how the vulnerability can be exploited, business impact, and potential solutions. This structure ensures that the responses obtained are well-organized and present the comprehensive information one would expect in a professional audit report, thereby laying the groundwork for LLMs to generate thorough audit reports in the future.
@@
C. Implications on the proposed approach on industrial use
@@
VI Data Availability
All the scripts and results that have been used during our experiments can be found in this zenodo repository https://zenodo.org/records/10530868
VII Related Works
@@
Static analysis involves examining the contract’s code without executing it. The goal is to identify vulnerabilities by analyzing the code’s structure, syntax, and logic. Tools such as Mythril, Manticore, and Slither are widely used for static analysis of Solidity code. Mythril, for example, uses static analysis, symbolic execution, and fuzzing for the detection of vulnerabilities. It also performs control flow and data flow analysis to understand how data moves through the contract. Its static analysis approach involves disassembling ethereum smart contract bytecode to an intermediate representation, typically leveraging an abstract syntax tree (AST). This analysis is augmented with pattern matching and heuristic techniques to identify known vulnerability patterns such as reentrancy, integer overflow and underflow, and other smart contract vulnerabilities. nevertheless, these tools produce a high ratio of false positive/negative. [12]
@@
Another approach involves using formal verification techniques [21] to mathematically prove the correctness of the smart contract. Formal verification involves using mathematical models to check the correctness of the code. This approach can be used to identify logical errors and ensure that the smart contract behaves as intended[22].
@@
Recently the application of machine learning and deep learning techniques to analyze solidity smart contracts has gained significant traction [23, 24]. Deep learning models, especially recurrent neural networks (RNNs) and their variant, Long Short-Term Memory (LSTM) networks, have been employed to capture the sequential nature of code and its inherent patterns [24]. LSTMs, with their ability to remember long-term dependencies, are particularly suited to recognizing intricate patterns in solidity code that might be indicative of vulnerabilities. While this approach have shown outstanding results, its generalization on unseen smart contracts written in a new version of solidity remains a concern. Also these models are not able to generate a human like explanations of the found vulnerabilities and the mitigation steps. On the other hand, GPT has attracted a lot of attention from the research community regarding its ability to code review and generation. Particularly in [10], a comprehensive evaluation of GPT3.5 and GPT-4 code interpreter for vulnerability detection in smart contract is illustrated. This work suggest to use prompt engineering for the auditing of smart contracts. In particular, it first uses GPT to generate the prompt that describes the task of auditing, then this prompt is used along with the full source code of the contracts as input to GPT. To measure the performance of GPT, the work considered only the Precision, Recall and F-measure metrics. however no qualitative evaluation of the generated output have been made. The work has not considered the non-deterministic nature of the GPT model regarding the prompt design. Also this work did not evaluate the capacity of GPT for the task of explaining the found vulnerabilities and the mitigation steps. The experiments have shown that the prompt length, i.e the contract length, can affect negatively the detection performance when exceeding the token limit.
@@
Similarly, in [9] a quantitative evaluation of the performance of GPT-4 code interpreter for the task vulnerability detection have been performed. Particularly, the prompt design included a description of a specific attack type besides to the whole source code of the contract. Common decentralized finance attacks have been classified into 38 Common weakness enumeration, CWE, and each smart contract have been prompted along with a description of the specific attack. Although the results of the experiment have shown great potential for GPT-4 code interpreter for detecting vulnerabilities, this experiment have shown that GPT performance is inversely proportional to contract length.
VIII Threat to validity
Internal Threat: A potential threat to the internal validity of our study arises from our automated script designed for extracting the Code Call List (CCL). Particularly when importing interfaces, complications may emerge. Due to various possible implementations for the same function signature, our algorithm might erroneously select an incorrect imported function. To alleviate this concern, we meticulously cross-examined all constructed CCLs, ensuring the correct implementation for each function was chosen.
In real world application, the smart contracts are often commented. The comments serves as a documentation of the written code. It also helps to enhance the code readability and clarity which provides the developers and auditors with deeper understanding of the aimed functionality of the code. For all these reasons, we did not remove the comments from the code. Our evaluation is performed on commented code. However, different smart contracts including those got using reverse engineering does not have comments. This may decrease the performance of the proposed approach.
Another aspect to consider, it the type of variables. Since the CCL is composed of a list of functions which can belong or not to the same contract, a special attention should be given to the variables being used inside the functions. Functions belonging to different smart contracts can be using the same name for state variables and GPT is unable to distinguish the difference between the two variables since it does not have access to the state variable declaration area. In future works, we will be prompting that section along with the functions list.
@@ VIII Threat to validity
External Threat: Our evaluation leverages data from sources such as Smart Bugs and the SolidiFy-benchmark. This data encompasses labeled vulnerabilities and has been open-source since 2020. Given that this data might have been incorporated into GPT-4 code interpreter’s training, it poses a potential external threat. To mitigate this, we created a new dataset composed of unlabeled functions from the evaluation contracts. However, a lingering uncertainty remains: these functions might have been labeled by other datasets, though we have no confirmation of such overlaps.
The non-deterministic nature of Generative Pre-trained Models (GPTs) poses significant challenges and limitations, particularly in contexts where consistency and reliability are paramount. This inherent unpredictability arises from the models’ training on diverse and extensive datasets, which leads to a stochastic approach in generating responses. While this can yield creative and varied outputs, it also results in a lack of consistency, making the model’s reliability questionable in critical applications.
Slither is used for code flattening and function call graph construction in the suggested approach. Slither supports one file flattening. We presume the created file and dependency management process are error-free. Complex dependencies might cause issues while flattening, resulting in unreachable code. All contracts Slither fails to flatten are discarded.
IX Conclusion
@@
In this paper, we introduced three context-augmentation approaches to improve smart contract co-auditing with Large Language Models (LLMs) and designed prompts around these strategies. The proposed approaches address challenges in handling long code, augmenting audit process knowledge, and focusing on specific vulnerabilities or weaknesses, thus enhancing LLM effectiveness in smart contract co-auditing. Our empirical study shows that the CCL-based chunking method consistently outperforms full code prompting. In the realm of prompt design, a nuanced comparison between the CAQ-based and CWE-based prompts indicates that while both exhibit comparable proficiency, the CAQ-based prompt is more efficient and showcases a broader vulnerability detection scope. Notably, the most effective strategy for auditing vulnerabilities in Solidity code using the GPT-4 code interpreter emerges as a mix of CCL-based chunking with the CAQ-based prompt.

### v2

@@ A. Affirming the Efficacy of Our CCL-Based Chunking Approach
To compare with the static analysis tools, we run multiple vulnerability detection tools on the smartBugs-curated dataset. The results are shown in table 3. The table compares the performance of the proposed CAQ+CCL (GPT + context augmenting) approach with well-known vulnerability detection tools: Slither[11], Mythril[30], Oyente [^7] https://github.com/enzymefinance/oyente, and Surya [^8] https://github.com/Consensys/surya. These tools represent traditional auditing methods, with Slither and Oyente focusing on static analysis, Mythril combining static and dynamic analysis with symbolic execution, and Surya providing static analysis along with code visualization. The table shows that CAQ+CCL outperforms all these tools in terms of the number of vulnerabilities detected, with a detection rate of 83.57%. In comparison, Slither detected 102 vulnerabilities with a rate of 49.27%, Mythril identified 100 vulnerabilities with a rate of 48.30%, Oyente detected 90 vulnerabilities with a rate of 43.47%, and Surya only identified 21 vulnerabilities with a rate of 10.14%. This highlights the superior performance of CAQ+CCL, not only in detecting more vulnerabilities but also in providing explanations of how these vulnerabilities can be exploited, offering more comprehensive support for auditors compared to the other tools.
Table 3: Comparison with Vulnerability Detection Tools
Tool Name Type Number of Detected Vulnerabilities Detection Rate (%) Ground Truth
CAQ+CCL approach GPT + context augmenting 173 83.57 207
Slither Static analysis 102 49.27 207
Mythril Static and Dynamic Analysis and symbolic execution 100 48.30 207
Oyente Static Analysis 90 43.47 207
Surya Static Analysis and Code Documentation 21 10.14 207
B. Adequacy of Our CAQ-Prompt in Practical Application
Our meticulously designed CAQ prompt is structured around four core components: judgment of vulnerabilities, explaining how the vulnerability can be exploited, business impact, and potential solutions. This structure ensures that the responses obtained are well-organized and present the comprehensive information one would expect in a professional audit report, thereby laying the groundwork for LLMs to generate thorough audit reports in the future.
@@ B. Adequacy of Our CAQ-Prompt in Practical Application
C. Interpretation of the effectiveness of the CCL based chunking technique
The proposed approach aims to provide an efficient prompting techniques that aims to reduce the prompt sizes specially regarding complex codebase while providing GPT with enough data in order to analyze effectively the targeted functions and localize vulnerabilities. As discussed by authors in [18], As the input length increases, the model has to manage a larger token numbers, which can lead to difficulties in maintaining coherence and relevance. This is particularly challenging in code-related tasks where precise understanding and execution of instructions are crucial. While models like GPT-4 have a high token limit, they still face challenges when approaching this limit. The need to chunk inputs or manage tokens effectively can result in loss of context or important details, impacting the quality of generated audits.
Besides, GPT4 model rely on the self-attention mechanism to understand the relationships between words and their contexts. Self-attention allows each token to ”attend” to all other tokens in the input sequence. However, this attention mechanism scales quadratically with the sequence length ($O(n^{2})$ complexity). For long sequences, the amount of computational resources and memory required increases exponentially, making it harder for the model to effectively capture relationships [32].
As discussed in the section 4.3.2, the finding 2 and 3 suggest that there is minimal disparity between the two prompting technique and thus the most effective approach is the combination of CCL-based chunking and the CAQ-based prompt. This can be explained by pattern recognition employed by GPT4. GPT-4 is trained on vast datasets and is designed to recognize patterns and generalize information. Whether using CCL-based or CAQ-based prompting to find vulnerabilities in a smart contract, the llm applies its understanding of common vulnerabilities and patterns it has learned for the training data.
The architecture of GPT models allows them to process and generate language based on learned patterns rather than specific rules. This means that for tasks like vulnerability detection, the model relies heavily on its training data’s statistical properties rather than a deterministic set of rules for pattern recognition [3].
D. Interpretation of the performance of GPT using CWE and CAQ prompt
The observed similarity in performance between CAQ prompts (generic prompts focusing on all vulnerabilities) and CWE prompts (specific prompts targeting particular vulnerabilities) can be attributed to the fundamental architecture and operation of large language models (LLMs). Transformer-based LLMs, such as OpenAI Codex, rely on identifying and applying patterns learned during pre-training to the input context. Generic prompts likely activate a broad range of the model’s knowledge, encompassing both general and specific vulnerability patterns. This aligns with findings by [4], which demonstrated Codex’s ability to handle diverse programming tasks, revealing that generic queries often leverage the model’s extensive knowledge base effectively.
The retrieval mechanisms of LLMs further explain this overlap. Models tend to retrieve contextual patterns based on the input query while inherently including related peripheral information due to the interconnected structure of their training data. Even when tasked with identifying specific vulnerabilities, the model’s retrieval process draws upon a broader set of associated patterns. This phenomenon has been highlighted in studies like [36], which explore the dense interconnections within LLM training data, facilitating the retrieval of comprehensive knowledge irrespective of prompt specificity.
Additionally, prompt optimization research sheds light on this behavior. Studies such as [5] indicate that carefully crafted prompts can enhance performance in tasks requiring deep reasoning or nuanced outputs. However, tasks that depend heavily on pre-trained knowledge, such as vulnerability detection, tend to exhibit minimal sensitivity to prompt specificity. Generic prompts already activate the broad retrieval patterns necessary to address most specialized queries, resulting in limited differences in performance between general and specific prompts.
These findings suggest that LLMs exhibit a degree of prompt invariance, particularly in domains where their pre-trained knowledge is expansive and well-integrated. For practical applications like smart contract auditing, this insight implies that focusing on robust generic prompts may suffice, reducing the need for extensive specialized prompt engineering. By leveraging the inherent strengths of LLMs, such an approach can streamline workflows while maintaining high task performance.
E. Implications on the proposed approach on industrial use
@@
6 Related Works
@@
Static analysis involves examining the contract’s code without executing it. The goal is to identify vulnerabilities by analyzing the code’s structure, syntax, and logic. Tools such as Mythril, Manticore[24], and Slither[11] are widely used for static analysis of Solidity code. Mythril, for example, uses static analysis, symbolic execution, and fuzzing for the detection of vulnerabilities. It also performs control flow and data flow analysis to understand how data moves through the contract. Its static analysis approach involves disassembling ethereum smart contract bytecode to an intermediate representation, typically leveraging an abstract syntax tree (AST). This analysis is augmented with pattern matching and heuristic techniques to identify known vulnerability patterns such as reentrancy, integer overflow and underflow, and other smart contract vulnerabilities. nevertheless, these tools produce a high ratio of false positive/negative .
@@
Another approach involves using formal verification and symbolic execution techniques [9, 33] to mathematically prove the correctness of the smart contract. Formal verification involves using mathematical models to check the correctness of the code. This approach can be used to identify logical errors and ensure that the smart contract behaves as intended[1].
@@
Recently, the application of machine learning and deep learning techniques to analyze Solidity smart contracts has gained significant traction [23, 28]. Deep learning models, especially Recurrent Neural Networks (RNNs) and their variant, Long Short-Term Memory (LSTM) networks, have been employed to capture the sequential nature of code and its inherent patterns [28]. LSTMs, with their ability to remember long-term dependencies, are particularly suited to recognizing intricate patterns in Solidity code that might be indicative of vulnerabilities. While this approach has shown outstanding results, its generalization to unseen smart contracts written in a new version of Solidity remains a concern. Also, these models are not able to generate human-like explanations of the found vulnerabilities and the mitigation steps. On the other hand, GPT has attracted a lot of attention from the research community regarding its ability to code review and generation [31, 2, 37, 21, 26]. Particularly in [3], a comprehensive evaluation of GPT-3.5 and GPT-4 code interpreter for vulnerability detection in smart contracts is illustrated. This work suggests using prompt engineering for the auditing of smart contracts. In particular, it first uses GPT to generate the prompt that describes the task of auditing; then, this prompt is used along with the full source code of the contracts as input to GPT. To measure the performance of GPT, the work considered only the Precision, Recall, and F-measure metrics. However, no qualitative evaluation of the generated output was made. The work did not consider the non-deterministic nature of the GPT model regarding the prompt design. Also, this work did not evaluate the capacity of GPT for the task of explaining the found vulnerabilities and the mitigation steps. The experiments showed that the prompt length, i.e., the contract length, can affect negatively the detection performance when exceeding the token limit.
@@
Similarly, in [6] a quantitative evaluation of the performance of GPT-4 code interpreter for the task of vulnerability detection has been performed. Particularly, the prompt design included a description of a specific attack type besides the whole source code of the contract. Common decentralized finance attacks have been classified into 38 Common Weakness Enumeration (CWE), and each smart contract has been prompted along with a description of the specific attack. Although the results of the experiment have shown great potential for GPT-4 code interpreter for detecting vulnerabilities, this experiment has shown that GPT performance is inversely proportional to contract length. The authors in [21] suggest combining fine-tuning and LLM-based agents to detect and explain vulnerabilities in smart contracts. It employs a multi-step process where the model first identifies potential issues, then reasons about the underlying causes, and finally ranks and critiques these explanations to ensure the most accurate and justified results.
In [33], authors focused on using static analysis techniques along symbolic execution for efficient vulnerability detection in smart contracts. By analyzing the inter-contract program dependency graph and performing symbolic execution on the smart contracts, they focus on the detection of reentrency vulnerabilities. While the evaluation of this approach have shown its efficiency for reentrency vulnerability detection, the work has not evaluated the approach on other types of smart contracts specific vulnerabilities. The authors in [34] present an innovative framework that utilizes a multi-agent conversational approach for detecting and analyzing vulnerabilities in smart contracts. This approach incorporates two primary strategies: Broad Analysis and Targeted Analysis. Although the evaluation demonstrates enhanced effectiveness in identifying vulnerabilities, the tool is highly resource-intensive.
Additionally, its performance can degrade when analyzing longer code segments. In [25], authors suggest an approach leveraging fine-tuned GPT-4 to identify and fix vulnerabilities in Solidity code by creating an annotated Control Flow Graph (CFG)
7 Threat to validity
Internal Threat: A potential threat to the internal validity of our study arises from our automated script designed for extracting the Code Call List (CCL). Particularly when importing interfaces, complications may emerge. Due to various possible implementations for the same function signature, our algorithm might erroneously select an incorrect imported function. To alleviate this concern, we meticulously cross-examined all constructed CCLs, ensuring the correct implementation for each function was chosen. In real-world applications, smart contracts are often commented. The comments serve as a documentation of the written code. It also helps to enhance the code readability and clarity, which provides the developers and auditors with a deeper understanding of the aimed functionality of the code. For all these reasons, we did not remove the comments from the code. Our evaluation is performed on commented code. However, different smart contracts, including those obtained through reverse engineering, may not have comments, which may decrease the performance of the proposed approach. Another aspect to consider is the type of variables. Since the CCL is composed of a list of functions which can belong or not to the same contract, a special attention should be given to the variables being used inside the functions. Functions belonging to different smart contracts can use the same name for state variables, and GPT is unable to distinguish the difference between the two variables since it does not have access to the state variable declaration area. In future works, we will be prompting that section along with the functions list.
@@ 7 Threat to validity
External Threat: Our evaluation leverages data from sources such as Smart Bugs and the SolidiFy-benchmark. This data encompasses labeled vulnerabilities and has been open-source since 2020. Given that this data might have been incorporated into GPT-4 code interpreter’s training, it poses a potential external threat. To mitigate this, we created a new dataset composed of unlabeled functions from the evaluation contracts. However, a lingering uncertainty remains: these functions might have been labeled by other datasets, though we have no confirmation of such overlaps. Dataset B mitigates the risk of model performance bias due to training overlap and enhances the evaluation’s real-world relevance by providing unseen smart contract code for auditing. The unseen smart contracts functions have been labeled by the two professional auditors from Quantstamp.
In our study, we define the divide the context into three components which are : code scoping, reporting scoping and assessment scoping. The context window is defined as the maximum number of tokens that an llm can process in a single prompt. As code length increases, the complexity of interactions between different parts of the code also increases. LLMs performance for the task of auditing smart contracts decrease especially for long code chunks. [18]. For this work, as the length of prompt is very important, we define the context of code as the code chunks and dependencies that are required for execution of the specific function we are auditing. While other works may have a different definition of the context, for this work we limit our definition of the context to the code chunks and dependencies a specific function requires to execute correctly The non-deterministic nature of Generative Pre-trained Models (GPTs) poses significant challenges and limitations, particularly in contexts where consistency and reliability are paramount. This inherent unpredictability arises from the models’ training on diverse and extensive datasets, which leads to a stochastic approach in generating responses. While this can yield creative and varied outputs, it also results in a lack of consistency, making the model’s reliability questionable in critical applications. Slither is used for code flattening and function call graph construction in the suggested approach. Slither supports one file flattening. We presume the created file and dependency management process are error-free. Complex dependencies might cause issues while flattening, resulting in unreachable code. All contracts that Slither fails to flatten are discarded.
8 Conclusion
@@
In this paper, we introduce three context-augmentation approaches to improve smart contract co-auditing with Large Language Models (LLMs) and design prompts around these strategies. The proposed approaches address challenges in handling long code, augmenting audit process knowledge, and focusing on specific vulnerabilities or weaknesses, thus enhancing LLM effectiveness in smart contract co-auditing. Our empirical study shows that the CCL-based chunking method consistently outperforms full code prompting. In the realm of prompt design, a nuanced comparison between the CAQ-based and CWE-based prompts indicates that while both exhibit comparable proficiency, the CAQ-based prompt is more efficient and showcases a broader vulnerability detection scope. Notably, the most effective strategy for auditing vulnerabilities in Solidity code using the GPT-4 code interpreter emerges as a mix of CCL-based chunking with the CAQ-based prompt.